### v1

# SMReferences ℓ

Abstract
We prove that for quantum spin chains with finite-range interactions, the existence of a specific conservation law known as the Reshetikhin condition implies the presence of infinitely many local conserved quantities, i.e., integrability. This shows that the entire hierarchy of conservation laws associated with solutions of the Yang–Baxter equation is already encoded in the lowest nontrivial conservation law. Combined with recent rigorous results on nonintegrability, our theorem strongly restricts the possibility of partially integrable systems that admit only a finite but large number of local conserved quantities. Our work establishes a rigorous foundation for the systematic identification of new integrable models and deepens the algebraic understanding of conservation-law structures in quantum spin chains.
Introduction
Quantum integrable models have long provided a unique window into the non‐perturbative behavior of quantum many‐body systems, from the exact solution of one‐dimensional quantum magnets [1] to the emergence of generalized hydrodynamics in far‐from‐equilibrium dynamics [2, 3]. Although a universally accepted definition of quantum integrability is lacking [4, 5], its essential feature is often regarded as the existence of an infinite hierarchy of local conserved quantities, which we adopt as the working definition in this Letter. Such conserved quantities encode much information about the initial state and constrain the dynamics in a way that generically prevents thermalization [6], enabling long-lived memory [7] and anomalous transport phenomena [8]. In contrast, generic quantum many-body systems admit only a few trivial conserved quantities, such as total energy and particle number. Therefore, integrability is a highly non-generic and fine-tuned property, and identifying integrable systems has been one of the central problems in rigorous statistical mechanics.
Traditionally, one establishes integrability by solving the system explicitly; for example, solving the Yang–Baxter equation and constructing a commuting family of local charges through the logarithmic derivative of the transfer matrix [9]. As an alternative, Reshetikhin proposed a simpler criterion [10]: the existence of a single third-order conservation law with a special form. Unlike approaches that rely on additional algebraic structures such as the Yang–Baxter equation or the self-duality [11], this criterion is algorithmically verifiable directly from the Hamiltonian. Since this Reshetikhin condition is a necessary condition for the $R$-matrix with difference form to solve the Yang–Baxter equation, it has long served as a guide in the search for new solvable systems [12, 13, 14, 15, 16, 17, 18]. Remarkably, in all known cases, systems satisfying the Reshetikhin condition have indeed turned out to be integrable. For this reason, it has been widely regarded as a heuristic test of integrability, as proposed by Grabowski and Mathieu [19].
Despite its wide use, however, it has remained an open question whether the Reshetikhin condition alone suffices to ensure integrability, namely the existence of infinitely many commuting local conserved quantities. It has often been argued that higher-order conservation laws and their mutual commutativity require additional algebraic input not captured by the third-order condition alone [14]. A recent preprint [20] claimed to resolve this question; however, its argument contains a critical flaw: a key equation employed in the proof is incorrect (see Supplemental Material [21] for details). In this Letter, we present the first complete and correct proof that any one‑dimensional quantum system satisfying the Reshetikhin condition necessarily gives rise to an infinite sequence of commuting local charges. Specifically, we show that the so-called boost operator [22, 23, 24] allows the iterative generation of higher‑order conservation laws. We further discuss a rigorous foundation for the empirical dichotomy of integrability: either a system possesses infinitely many local conserved quantities, or a few simple at most two-body conserved charges.
Setup and main result
We consider a quantum spin system on an infinite open chain, with sites labeled by $\mathbb{Z}$. Each site $i\in\mathbb{Z}$ has a local Hilbert space $\mathcal{H}_{i}\cong\mathbb{C}^{d}$, where $d<\infty$. We denote by $\mathcal{B}_{i}$ the set of bounded linear operators on $\mathcal{H}_{i}$, and by $\mathcal{B}_{0i}$ the subset of traceless operators. We call an operator at most $n$-local if it can be expressed as a sum of terms acting nontrivially on at most $n$ contiguous sites, and $n$-local if it is not (n1)(n-1)-local.
The Hamiltonian consists of nearest-neighbor interactions and on-site potentials:
$\hat{H}=\sum_{i\in\mathbb{Z}}\hat{h}^{(2)}_{i}+\sum_{i\in\mathbb{Z}}\hat{h}^{(1)}_{i}$, (1)
with $\hat{h}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0,i+1}$ and $\hat{h}^{(1)}_{i}\in\mathcal{B}_{0i}$. We assume Hermiticity and translational invariance. The mathematically rigorous definitions of the formal sum in Eq. (1) and of the commutators below are given in Supplemental Material [21].
A key ingredient in our analysis is the boost operator [22, 23, 24].
Definition 1.
Fix an arbitrary translationally invariant one-local operator $\hat{C}=\sum_{i\in\mathbb{Z}}\hat{c}_{i}$. The boost operator associated with the Hamiltonian $\hat{H}$ is defined by
$\hat{B}_{\hat{H}}\coloneqq\sum_{j\in\mathbb{Z}}j(\hat{h}^{(2)}_{j}+\hat{h}^{(1)}_{j})+\hat{C}$. (2)
We set $\hat{Q}^{(2)}\coloneqq\hat{H}$ and recursively $\hat{Q}^{(n+1)}\coloneqq[\hat{B}_{\hat{H}},\hat{Q}^{(n)}]$. By construction, $\hat{Q}^{(n)}$ is at most $n$-local. Although the full structure of $\hat{Q}^{(n)}$ is complicated, its $n$-local component admits the compact expression
$(-1)^{n}(n-2)!\sum_{i\in\mathbb{Z}}[\dots[[\hat{h}_{i}^{(2)},\hat{h}_{i+1}^{(2)}],\hat{h}_{i+2}^{(2)}],\dots,\hat{h}_{i+n-2}^{(2)}],$ (3)
which is nonzero whenever the injectivity condition (Eq. (11)) holds [25].
For any translationally invariant operator $\hat{Q}$ commuting with $\hat{H}$, the commutator $\hat{B}_{\hat{H}},\hat{Q}$] is again translationally invariant [26, 27]. Indeed, denoting by $\mathcal{T}:\mathcal{B}_{i}\to\mathcal{B}_{i-1}$ the translation operator, one finds $\mathcal{T}(\hat{B}_{\hat{H}})=\hat{B}_{\hat{H}}+\hat{H}$, and hence
$\mathcal{T}[\hat{B}_{\hat{H}},\hat{Q}]=[\mathcal{T}(\hat{B}_{\hat{H}}),\hat{Q}]=[\hat{B}_{\hat{H}}+\hat{H},\hat{Q}]=[\hat{B}_{\hat{H}},\hat{Q}$]. (4)
In particular, we have
$\hat{Q}^{(3)}$ $=[\hat{B}_{\hat{H}},\hat{H}$]
$=-\sum_{i\in\mathbb{Z}}[\hat{h}^{(2)}_{i},\hat{h}^{(2)}_{i+1}+\hat{h}^{(1)}_{i+1}+\hat{c}^{(1)}_{i}+\hat{c}^{(1)}_{i+1}$]
$\hskip 28.45274pt-\sum_{i\in\mathbb{Z}}[\hat{h}^{(1)}_{i},\hat{c}^{(1)}_{i}$]
$=-\sum_{i\in\mathbb{Z}}[\hat{h}_{\hat{C},i},\hat{h}_{\hat{C},i+1}],$ (5)
where $\hat{h}_{\hat{C},i}\coloneqq\hat{h}^{(2)}_{i}+\hat{h}^{(1)}_{i}+\hat{c}^{(1)}_{i}-\hat{c}^{(1)}_{i+1}\in\mathcal{B}_{i}\otimes\mathcal{B}_{i+1}$.
We now state our main result.
Theorem 2.
If $\hat{Q}^{(3)}$ is a conserved quantity of $\hat{H}$, i.e., $\hat{H},\hat{Q}^{(3)}]=0$, then for any $n\geq 2$, $\hat{Q}^{(n)}$ is a translationally invariant conserved quantity of $\hat{H}$. Moreover, the operators in $\{\hat{Q}^{(n)}\}_{n=2}^{\infty}$ mutually commute.
As noted above, for generic Hamiltonians the operator $\hat{Q}^{(n)}$ is $n$-local, so that each element of $\{\hat{Q}^{(n)}\}_{n=2}^{\infty}$ yields a distinct and nontrivial conserved quantity. Therefore, Theorem 2 demonstrates that a single conservation law, $\hat{Q}^{(3)}$, generates an infinite hierarchy of local conserved quantities, i.e., integrability. Since the condition $\hat{H},\hat{Q}^{(3)}]=0$, or its local version, is precisely the Reshetikhin condition [10], Theorem 2 can be summarized as the Reshetikhin condition implies integrability. This is equivalent to the positive resolution of Conjecture 2 of Grabowski and Mathieu [19] for the standard boost operator $\hat{B}_{\hat{H}}$ (the case of general boosts will be addressed later). Our result establishes, for the first time, that the third-order conservation law alone is sufficient to guarantee full quantum integrability.
Since commutativity of translationally invariant local operators is a local property (see Theorem S10 and Corollary S13 in Supplemental Material [21]), Theorem 2 extends to sufficiently large finite systems. Concretely, for a translationally invariant local operator $\hat{Q}=\sum_{i\in\mathbb{Z}}\hat{q}_{i}$ in the infinite system, we can define its finite-size counterpart $\hat{Q}_{[N]}=\sum_{i=1}^{N}\hat{q}_{i}$ on a chain of length $N$ with the periodic boundary condition. Then we have the following corollary.
Corollary 3.
If $\hat{Q}^{(3)}_{[N_{0}]}$ is a conserved quantity of $\hat{H}_{[N_{0}]}$, i.e., $\hat{H}_{[N_{0}]},\hat{Q}^{(3)}_{[N_{0}]}]=0$ for some $N_{0}\geq 6$, then $\hat{Q}^{(n)}_{[N]},\hat{Q}^{(m)}_{[N]}]=0$ for all $N\geq 4$ and $(n,m)$ satisfying $2\leq n,m\leq\frac{N}{2}$. In particular, $\hat{Q}^{(n)}_{[N]}$ is a conserved quantity of $\hat{H}_{[N]}$ for all $2\leq n\leq\frac{N}{2}$.
Proof of Theorem 2
If all $\hat{Q}^{(n)}$ are conserved quantities of $\hat{H}$, their mutual commutativity follows immediately from the Jacobi identity. In fact, the following standard lemma [15, 20] holds.
Lemma 4.
Let $M\geq 1$. If $\hat{H},\hat{Q}^{(2m+1)}]=0$ holds for all $1\leq m\leq M$, then $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ holds for $k+l\leq 2M+2$.
Proof.
We proceed by induction on $M$. For $M=1$, the only nontrivial case is $\hat{H},\hat{Q}^{(4)}]=0$. This can be shown from the Jacobi identity as follows:
$[\hat{H},\hat{Q}^{(4)}]$ $=[\hat{H},[\hat{B}_{\hat{H}},\hat{Q}^{(3)}]]$
$=[\hat{Q}^{(3)},[\hat{B}_{\hat{H}},\hat{H}]]+[\hat{B}_{\hat{H}},[\hat{H},\hat{Q}^{(3)}]]$
$=[\hat{Q}^{(3)},\hat{Q}^{(3)}]=0$ (6)
Assume that the claim holds up to $M=M^{\prime}-1\geq 1$, and consider $M=M^{\prime}$. The only nontrivial part to prove is that $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ for $2M^{\prime}+1\leq k+l\leq 2M^{\prime}+2$. First, consider the case $k+l=2M^{\prime}+1$. By the Jacobi identity, all such commutators are proportional to each other:
$[\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime}+1)}]$ $=-[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime})}]+[\hat{B}_{\hat{H}},[\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime})}]]$
$=-[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime})}]$
$=[\hat{Q}^{(4)},\hat{Q}^{(2M^{\prime}-1)}]-[\hat{B}_{\hat{H}},[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime}-1)}]]$
$=[\hat{Q}^{(4)},\hat{Q}^{(2M^{\prime}-1)}]$
$=\dots$
$=(-1)^{M^{\prime}-1}[\hat{Q}^{(M^{\prime}+1)},\hat{Q}^{(M^{\prime}+2)}]$ (7)
where we used the induction hypothesis $\hat{Q}^{(k+1)},\hat{Q}^{(2M^{\prime}-k+1)}]=0$. In particular, for $k=1$ the hypothesis already gives $\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime}+1)}]=0$. Hence, $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ holds for all $k+l=2M^{\prime}+1$.
For $k+l=2M^{\prime}+2$, the case $k=1$ follows similarly as in Eq. (6). Moreover, proportionality as in Eq. (7) again holds, so $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ is established for all pairs $(k,l)$ with $k+l=2M^{\prime}+2$. ∎
From this lemma, it suffices to prove Theorem 2 by showing: If $\hat{H},\hat{Q}^{(n)}]=0$ holds for some $n\geq 3$, then $\hat{H},\hat{Q}^{(n+1)}]=0$ also holds. The key step is the following observation.
Lemma 5.
If $\hat{H},\hat{Q}^{(3)}]=0$ and $\hat{H},\hat{A}]=0$ holds for some local operator $\hat{A}$, then $\hat{H},[\hat{H},[\hat{B}_{\hat{H}},\hat{A}]]]=0$.
Proof.
From the Jacobi identity, we have
$[\hat{H},[\hat{B}_{\hat{H}},\hat{A}$]] $=-[\hat{Q}^{(3)},\hat{A}]+[\hat{B}_{\hat{H}},[\hat{H},\hat{A}$]]
$=-[\hat{Q}^{(3)},\hat{A}$]. (8)
Since both $\hat{Q}^{(3)}$ and $\hat{A}$ commute with $\hat{H}$, we obtain
$[\hat{H},[\hat{H},[\hat{B}_{\hat{H}},\hat{A}$]]] $=-[\hat{H},[\hat{Q}^{(3)},\hat{A}]]=0,$ (9)
which proves the claim. ∎
In view of this lemma, it remains to establish the following.
Lemma 6.
For a translationally invariant $n(\geq 1)$-local operator $\hat{A}^{(n)}$, if $\hat{H},[\hat{H},\hat{A}^{(n)}]]=0$, then $\hat{H},\hat{A}^{(n)}]=0$.
Proof.
Consider a finite system of size $N$ with the periodic boundary condition. Note that the commutativity of translationally invariant local operators is a local property, and hence independent of $N$. Thus, for sufficiently large $N$ compared with $n$, the assumption can be replaced with the finite-size relation $\hat{H}_{[N]},[\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}]]=0$ (See Corollary S13 of Supplemental Material [21] for a formal proof). By considering the Hilbert-Schmidt norm of $\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}$], we obtain
[…]
$=\operatorname{Tr}({\hat{A}^{(n)\dagger}_{[N]}}[\hat{H}_{[N]},[\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}]])=0,$ (10)
which implies $\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}]=0$. Therefore, the corresponding infinite-system relation $\hat{H},\hat{A}^{(n)}]=0$ also holds. ∎
Together, these lemmas establish Theorem 2. It should be noted that we cannot dispense with the assumption of translational invariance in Lemma 6, since $\hat{H},[\hat{H},\hat{B}_{\hat{H}}]]=0$ does not necessarily imply that $\hat{Q}^{(3)}=-[\hat{H},\hat{B}_{\hat{H}}]=0$.
Dichotomy on integrability
For quantum spin chains with local interactions, a generic system that is not integrable is expected to be (completely) nonintegrable, meaning that it possesses no local conserved quantities other than trivial ones such as the Hamiltonian and the total magnetization. Indeed, this expectation has been rigorously confirmed for spin-$1/2$ chains with parity-symmetric interactions [28, 29, 30] by applying the method developed in Ref. [31] to rigorously prove nonintegrability for specific systems. However, these results rely on an ad hoc, case-by-case verification of nonintegrability.
In the following, we address how to rigorously establish this belief for a broader class of Hamiltonians, namely the dichotomy on integrability: either there exist infinitely many local conserved quantities, or there exist none beyond the trivial ones. Specifically, we combine Theorem 2, which provides a sufficient condition for integrability, with a sufficient condition for nonintegrability that the author has recently proved [25].
We assume that the interaction $\hat{h}^{(2)}$ is sufficiently complicated in the sense of Ref. [25]. First, for any $\hat{a}^{(1)}\in\mathcal{B}_{0}$, we assume that
$=0$ $\;\Rightarrow\;\hat{a}^{(1)}=0,$ (11)
$[\hat{a}^{(1)}\otimes\hat{I},\hat{h}^{(2)}_{1}]=0$ $\;\Rightarrow\;\hat{a}^{(1)}=0,$
which is a slightly stronger version of the condition that an on-site operator $\hat{a}^{(1)}$ is not a conserved quantity.
Next, for a family of operators $\hat{a}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0i+1}$, we assume that
$\hat{h}_{i}^{(2)},\hat{a}^{(2)}_{i+1}]=[\hat{a}^{(2)}_{i},\hat{h}_{i+1}^{(2)}]\quad(\forall i)\;\Rightarrow\;\hat{a}^{(2)}_{i}=\alpha\,\hat{h}_{i}^{(2)}\quad(\forall i),$ (12)
where $\alpha\in\mathbb{C}$ is a constant independent of $i$. This condition implies that the only two-local conserved quantity is the Hamiltonian itself.
Proposition 7.
Let $\mathbb{H}$ be the set of Hamiltonians satisfying assumptions (11) and (12). For any $\hat{H}\in\mathbb{H}$, suppose that the following holds.
  1. (A)
    Let $\hat{Q}^{(3)}=[\hat{B}_{\hat{H}},\hat{H}$]. If there exists a two-local operator $\hat{R}^{(2)}$ such that $\hat{H},\hat{Q}^{(3)}+\hat{R}^{(2)}]$ is at most two-local, then $\hat{H},\hat{Q}^{(3)}]=0$.
Then exactly one of the following holds.
  • $\hat{H}$ is integrable in the sense that there exist mutually commuting $k$-local conserved quantities for all $k\geq 2$.
  • $\hat{H}$ is nonintegrable in the sense that the only $k$-local conserved quantities are the Hamiltonian itself ($k=2$) and one-local operators.
Proof.
From condition (A), one of the following holds.
  • $\hat{H}$ commutes with $\hat{Q}^{(3)}$.
  • For any two-local operator $\hat{R}^{(2)}$, the commutator $\hat{H},\hat{Q}^{(3)}+\hat{R}^{(2)}]$ is three-local.
In the former case, Theorem 2 implies the first conclusion. In the latter case, the result of Ref. [25] implies the second. ∎
Therefore, once condition (A) is established, the system admits a complete dichotomy: it is either integrable or nonintegrable. Conversely, if a “partially integrable” system were to exist, then within the framework of assumptions (11) and (12), it would necessarily provide a counterexample to condition (A).
Proof of the Grabowski–Mathieu conjecture for general boosts
Grabowski and Mathieu conjectured that the existence of a boost operator $\hat{B}$ with $\hat{B},\hat{H}$] conserved implies integrability of $\hat{H}$ [19]. Theorem 2 established this for the canonical boost $\hat{B}=\hat{B}_{\hat{H}}$ which is obtained from the Yang–Baxter equation. The same reasoning applies to general boosts for Hamiltonians satisfying assumptions (11) and (12). Assume that $\hat{B}$ is Hermitian and of the form
$\hat{B}=\sum_{j\in\mathbb{Z}}\hat{b}^{(2)}_{j}+\sum_{j\in\mathbb{Z}}\hat{b}^{(1)}_{j}$, (13)
with $\hat{b}^{(2)}_{j}\in\mathcal{B}_{0j}\otimes\mathcal{B}_{0j+1}$ and $\hat{b}^{(1)}_{j}\in\mathcal{B}_{0j}$. That is, $\hat{B}$ is at most two-local. Setting $\hat{Q}^{(2)}\coloneqq\hat{H}$ and $\hat{Q}^{(n+1)}=[\hat{B},\hat{Q}^{(n)}]$, we obtain the following theorem.
Theorem 8.
Let $\hat{H}$ satisfy assumptions (11) and (12). If $\hat{Q}^{(3)}=[\hat{B},\hat{H}$] is a nonzero conserved quantity, then for any $n\geq 2$, $\hat{Q}^{(n)}$ is a translationally invariant conserved quantity of $\hat{H}$. Moreover, the operators in $\{\hat{Q}^{(n)}\}_{n=2}^{\infty}$ mutually commute. In addition, $\hat{B}$ is restricted to the form
$\hat{B}\propto\sum_{j\in\mathbb{Z}}j(\hat{h}^{(2)}_{j}+\hat{h}^{(1)}_{j})+\hat{R}+a\hat{H}$, (14)
with $a\in\mathbb{R}$, where $\hat{R}$ is one-local and satisfies
$\hat{H},\mathcal{T}(\hat{R})-\hat{R}]=0.$ (15)
Proof.
The proof that $\hat{B}$ must take the form of Eq. (14) is given in End Matter. Accepting this fact, we may, without loss of generality, write $\hat{B}=\hat{B}_{\hat{H}}+\hat{R}$.
Since Lemmas 4, 5 (with $\hat{B}_{\hat{H}}$ replaced by $\hat{B}$), and 6 remain valid, it suffices to show $\hat{Q}^{(n)}$ is translationally invariant. If $\hat{Q}^{(k)}$ is translationally invariant, then
$\mathcal{T}(\hat{Q}^{(k+1)})-\hat{Q}^{(k+1)}$ $=[\mathcal{T}(\hat{B})-\hat{B},\hat{Q}^{(k)}]$
$=[\hat{H}+\hat{X},\hat{Q}^{(k)}]=[\hat{X},\hat{Q}^{(k)}],$ (16)
where $\hat{X}\coloneqq\mathcal{T}(\hat{R})-\hat{R}$. For $k=2$ this vanishes, proving translational invariance of $\hat{Q}^{(3)}$. For general $k$, the one-locality of $\hat{X}$ again ensures $\hat{X},\hat{Q}^{(k)}]=0$ (see Proposition 12 in End Matter), and hence $\hat{Q}^{(k+1)}$ is also translationally invariant. Thus, all $\hat{Q}^{(n)}$ are translationally invariant, and commutativity follows as in Theorem 2. ∎
Theorem 8, which does not assume any specific form of $\hat{B}$, provides a complete proof of the Grabowski–Mathieu conjecture (Conjecture 2 in Ref. [19]) for Hamiltonians satisfying assumptions (11) and (12): the existence of a boost generating a single nonzero local conserved quantity implies integrability. Eq. (14) further shows that any boost coincides with the standard $\hat{B}_{\hat{H}}$ up to a one-local term.
In contrast, in systems not satisfying assumption (12), such as the XY model, there exists a boost operator different from the standard $\hat{B}_{\hat{H}}$, which in turn generates two independent families of local conserved quantities [32]. This also follows from the following proposition, which generalizes Theorem 2.
Proposition 9.
Suppose that $\hat{Q}^{(3)}$ is a conserved quantity of $\hat{H}$ and that $\hat{X}=\mathcal{T}(\hat{B})-\hat{B}$ is translationally invariant. If
$\hat{X},[\hat{B},\hat{X}]]=0,$ (17)
then all operators in $\{\hat{Q}^{(n)}\}_{n=2}^{\infty}$ are translationally invariant, mutually commuting, and commute with $\hat{X}$.
The proof is essentially the same as that of Theorem 2, except that commutativity with $\hat{X}$ must also be shown (see End Matter).
Generalization to parametrized Hamiltonians
Certain integrable models, such as the Hubbard model, have an $R$-matrix that cannot be expressed as a function of a single real variable. As a result, they do not satisfy the standard Reshetikhin condition and, moreover, they admit no boost operator [32]. Instead, they are associated with families of parametrized Hamiltonians obeying a generalized Reshetikhin condition [33]. As a counterpart to Theorem 2, we show that if a parametrized Hamiltonian satisfies this generalized condition, then it is integrable.
We assume that the Hamiltonian takes the form
$\hat{H}(\lambda)=\sum_{i\in\mathbb{Z}}\hat{h}^{(2)}_{i}(\lambda)+\sum_{i\in\mathbb{Z}}\hat{h}^{(1)}_{i}(\lambda),$ (18)
where $\lambda$ belongs to some open subset $U\subset\mathbb{R}^{D}$. For an arbitrary smooth $v:U\to\mathbb{R}^{D}$, define
$\hat{Q}^{(2)}(\lambda)$ $\coloneqq\hat{H}(\lambda),$ (19)
$\hat{Q}^{(n+1)}(\lambda)$ $\coloneqq[\hat{B}_{\hat{H}(\lambda)},\hat{Q}^{(n)}(\lambda)]+(v\cdot\nabla)\hat{Q}^{(n)}(\lambda),$ (20)
where $v\cdot\nabla=\sum_{a=1}^{D}v_{a}\partial_{\lambda_{a}}$. In particular,
$\hat{Q}^{(3)}(\lambda)=-\sum_{i\in\mathbb{Z}}[\hat{h}_{i}(\lambda),\hat{h}_{i+1}(\lambda)]+(v\cdot\nabla)\hat{H}(\lambda),$ (21)
where $\hat{h}_{i}(\lambda)=\hat{h}^{(2)}_{i}(\lambda)+\hat{h}^{(1)}_{i}(\lambda)+\hat{c}^{(1)}_{i}(\lambda)-\hat{c}^{(1)}_{i+1}(\lambda)$ for some translationally invariant $1$-local operator $\hat{C}(\lambda)$. The generalized Reshetikhin condition requires $\hat{Q}^{(3)}(\lambda),\hat{H}(\lambda)]=0$ for all $\lambda\in U$, or its local form [33].
In parallel with the standard case (Theorem 2), the following holds.
Theorem 10.
If $\hat{Q}^{(3)}(\lambda)$ is a conserved quantity of $\hat{H}(\lambda)$ for every $\lambda\in U$, then $\hat{Q}^{(n)}(\lambda)$ is a translationally invariant conserved quantity of $\hat{H}(\lambda)$ for all $n\geq 2$,